OpenStudy (anonymous):

Simplify the following expression. 3x − 6y + 2z − 5x − 3y + 8z −2x + 9y + 10z −2x + 9y − 10z −2x − 9y + 10z 2x − 9y + 10z

OpenStudy (anonymous):

3x-5x=-2x -6y-3y=-9y 2z+8z=10z so add them

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

your welcome ! :D

OpenStudy (anonymous):

;)